Suppose the yearly inflation rate from 2014 to 2020 is 15%, the table that costs $800 at the start of 2014 costs $920 at the start of 2020, and so on. What equation represents the cost of the table from the year 2014 to 2020?
Answer by ikleyn(53750)

.
Suppose the yearly inflation rate from 2014 to 2020 is 15%, the table that costs $800 at the start of 2014
costs $920 at the start of 2020, and so on. What equation represents the cost of the table from the year 2014 to 2020?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

    Cost(2020) = Cost(2014)*(1+0.15) = 800*1.15.    ANSWER

Solved.